OpenStudy (anonymous):

Hitomi has a bag with four chips in it numbered 1 to 4. She randomly draws a chip and replaces it three times. What is the probability that Hitomi will draw a 1 all three times?

OpenStudy (anonymous):

The probability of drawing a 1 once is 1/4, since 1 is one 1 out of 4 tiles. To find it four times, your equation will look like this \[\frac{ 1 }{ 4 }\times \frac{ 1 }{ 4}\times \frac{ 1 }{ 4}\times \frac{ 1 }{ 4}\] Since she replaces it each time, the denominator stays constant.